OpenStudy (anonymous):

Kelsey painted one-fourth of her bedroom in three-sevenths of an hour. At this rate, how long would it take her to paint the entire room? three over twenty-eight of an hour twenty-eight over three of an hour twelve over seven of an hour seven over twelve of an hour

OpenStudy (anonymous):

@Great_And_Powerful

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

is it 12 over 7

OpenStudy (igreen):

We have to divide: \(\dfrac{3}{7} \div \dfrac{1}{4}\) Or multiply: \(\dfrac{3}{7} \times \dfrac{4}{1} \rightarrow \dfrac{3 \times 4}{7 \times 1}\) Simplify

OpenStudy (igreen):

Yep, you got it.. \(\dfrac{3}{7} \times \dfrac{4}{1} \rightarrow \dfrac{3 \times 4}{7 \times 1} \rightarrow \dfrac{12}{7}\)

OpenStudy (anonymous):

thaks

OpenStudy (anonymous):

Margie calculated that she would spend $175 on school supplies this year. She actually spent $97.50 on school supplies. What is Margie’s percent of error? 77.3% 79.5% 44.2% 4.4%

OpenStudy (anonymous):

and is it c

OpenStudy (igreen):

One question per post, close this one and open a new one.